OpenStudy (anonymous):

[4 - (3c - 1)][6 - ( 3c - 1)]

OpenStudy (anonymous):

[4 - 1(3c - 1)][6 - 1( 3c - 1)] Distribute the -1 to 3c-1 in both to get : (4-3c+1)(6-3c+1) add like terms in the () and get (-3c+5)(-3c+7) and FOIL.

OpenStudy (anonymous):

a - sign in an equation for example 8-7 is actually 8+ (-7) which is what I did with the - in the above equation.

OpenStudy (anonymous):

It had an invisible 1 so I took out the one and made it [4 + (- 1)(3c - 1)][6 + (- 1)( 3c - 1)], both are the same thing.

OpenStudy (anonymous):

so i foiled it and i and did as you said but it came out as incorrect

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

27c+35

OpenStudy (anonymous):

Nooooo That's wrong. Can you post up the steps you took to foil so I can see what you did?

OpenStudy (anonymous):

3xy X 4xy then 3xy X 2 then -1 X 4xy then -1 X2 and i got 12xy +6xy-4xy-2

OpenStudy (anonymous):

nevermind i got 14xy-2

OpenStudy (anonymous):

Uhm...? Are you still talking about this problem..?

OpenStudy (anonymous):

For this problem you should have got : (-3c+5)(-3c+7) 9c^2 -21c - 15c + 35 9c^2 - 36c + 35

OpenStudy (anonymous):

And you have to remember to multiply the variables too. Not just the numbers.

OpenStudy (anonymous):

A variable multiplied by a variable = x * x = x^2, or xy * xy = x^2y^2, or even x * y = xy.

OpenStudy (anonymous):

omg am so sorry i jst get lost its too many math problems that i dont even get

OpenStudy (anonymous):

Lol. It's okay.

OpenStudy (anonymous):

but am having the same problems with this one as well (2x-3y)(4x-y)

OpenStudy (anonymous):

That one is 2x * 4x = 8x^2, 2x * y = 2xy, -3y * 4x = -12xy, and last -3y * -y = 3y^2, put it together and you will get 8x^2 + 2xy - 12xy + 3y^2 add like terms, the final solution should be 8x^2 - 10xy + 3y^2